# Diving into the Fusion of Monocular Priors for Generalized Stereo Matching

### v1

@@ 4.4 Ablation Study and Analysis
The Effectiveness of Each Module. As shown in Table 5, the baseline model performs better without context features, indicating that monocular priors are susceptible to domain bias when data is limited. By incorporating less-biased monocular priors from a pre-trained large monocular network in the monocular encoder (ME), generalization performance is significantly improved, highlighting the importance of robust monocular priors in the wild world. Comparing the Baseline + ME with Baseline in Table 1, its performance becomes worse than Raft-Stereo, showing that it is easy to suffer over-confidence when simply fusing monocular features with disparities during iterative disparity update. The iterative direct fusion method (IDF) fuses monocular depth and binocular disparity through direct concatenation and convolution at each iteration. Compared to this approach, our iterative local fusion (ILF) is more robust to noise in binocular disparity, resulting in superior performance. The post-fusion method (PF) fuses monocular depth with the optimized binocular disparity from the previous iteration without registration. Compared to this approach, our global fusion (GF) achieves better compatibility between monocular depth and binocular disparity, mitigating the noise caused by scale ambiguity during fusion. Our iterative local fusion and global fusion modules further enhance performance and improve model robustness when combined with the monocular encoder. It is also noted that the time cost of the Baseline is 0.32s while our model is 0.4s. Even though it involves a VTF model, benefiting from the elegant and controllable design, our model barely raises the time cost.
The Analysis of Iterative Local Fusion. We also analyze the specific configurations of iterative local fusion. As shown in Table 6, the fixed weights in the LBP-like operation have a slight impact on performance, with a kernel size of ${3,5}$ providing optimal results. We also try to use convolutions with learnable weights to replace LBP-like convolutions. Comparing L(4), L(9-10) with L(6-7), we find that fixed-weight convolutions are more robust than learnable convolutions, and deeper learnable convolutions produce worse results. This is because limited data makes monocular-related learning unreliable for generalization, whereas manually designed convolutions incorporate prior knowledge and are less affected by data bias. Using a sigmoid function after LBP-like convolutions further improves overall performance. The amplitude parameter does not show a significant influence. We also visualize the local ordering map in Figure 3. The local ordering maps of predicted disparity gradually become similar to the result of monocular depth as the iteration increases. For more visualizations, please refer to our supplemental materials.
@@ 5 Conclusion
Appendix A Intuition behind Monocular Depth Model
@@
Method Additional Data/Aug Booster
@@
Method Additional Data/Aug Booster
@@
Appendix B More Results on Booster
We provide additional results on the Booster dataset across various material types. From class 00 to $3$, the materials become increasingly transparent and/or specular. As shown in Tables 8 and 9, our method outperforms state-of-the-art approaches on transparent and/or specular objects (classes $1$ to $3$), while achieving comparable results in normal regions (class 00). The normal regions of the Booster dataset mainly consist of regular objects, flat surfaces, or highly textured areas. Consequently, NerfStereo, which incorporates additional stereo data, performs particularly well in these regions. This indicates that stereo matching effectively captures fine-grained details, whereas monocular depth estimation excels in perceiving coarse shapes. As illustrated in Figures 6 and 7, binocular disparity provides greater detail compared to monocular depth. Our method disentangles monocular depth and binocular disparity, allowing model to leverage both monocular and stereo data, and explore the fusion of monocular priors effectively.
750$\times$2484 1125$\times$3726 1500$\times$4968 1688$\times$5589 1875$\times$6210
@@ Appendix B More Results on Booster
Table 10: Memory comparison. We evaluate the memory consumption of each model, excluding the feature encoder module, to ensure a fair comparison of backbones during inference. The evaluation is performed across different resolutions. ’reg’ denotes pre-computation of the entire cost volume, allowing for look-up operations at each iteration, while ’alt’ refers to dynamically computing a thin cost volume at each iteration. The 384 and 640 represent the maximum disparity range used for the resolution of 750$\times$2484. ’-’ indicates that the GPU does not support the model at the given resolution.
@@
Baseline + FE-DepthAnything 3.26$\pm$0.03 28.73$\pm$0.28
Baseline + FE-MASt3R 4.41$\pm$0.40 26.83$\pm$0.57
Baseline + ME + ILF + GF 1.15$\pm$0.01 8.35$\pm$0.04
Table 11: The effectiveness of each module. Baseline is RAFTStereo, while ME is our monocular encoder, ILF is our iterative local fusion, and GF is our global fusion. FE-DepthAnything means using DepthANything v2 to replace the original feature extractor. FE-MASt3R means using MASt3R to replace the original feature extractor.
Appendix C More Analysis about Memory
@@
Appendix D More Visualization for Generalized Stereo Matching
@@
We provide additional visualizations of generalized stereo matching in Figures 8, 9, 10, 11, and 12. The visualizations span a variety of environments, ranging from open outdoor scenes (e.g., driving scenarios), to semi-open outdoor scenes (e.g., playgrounds), and to enclosed indoor scenes (e.g., rooms, tables). The results demonstrate that our method generalizes effectively to the wild world, achieving strong performance even when trained only on a limited amount of synthetic stereo data.
Appendix E Ablation Study
E.1 More Analysis of Backbone
@@
In addition to replacing the context network with the pre-trained DepthAnything v2 [49], we also experimented with replacing the feature extractor for cost volume construction using DepthAnything v2 [49] and MASt3R [mast3r_arxiv24, dust3r_cvpr24]. As shown in Table 11, the results become worse after replacing the feature extractor for cost volume construction with DepthAnything v2 or MASt3R. Moreover, there is a bug with the A40 GPU that causes memory issues when converting the alternate correlation function from dot product to Euclidean distance during training. Therefore, the model with MASt3R was trained using the original correlation function with dot product, where additional learnable convolution layers are further used after MASt3R for feature extraction.
E.2 More Analysis of Iterative Local Fusion
@@
We provide additional visualizations of the intermediate results from the iterative local fusion process in Figures 13, 14, 15, 16, 17, 18, 19, and 20. As the iterations progress, the ordering maps generated from binocular disparity gradually become smoother. The convolution layers learn the differences between ordering maps generated from binocular disparity and monocular depth, allowing the guidance to focus more effectively on non-smooth regions, thereby significantly affecting disparity update.
E.3 More Analysis of Components in Global Fusion
@@
We present more visualization for the intermediate results of global fusion in Figure 13, 14, 15, 16, 17, 18, 19, and 20. The visualization shows that the registration of monocular depth is different for each pixel, particularly on different objects. Since the monocular depth from DepthAnything is scale ambiguity but not absolute depth before registration, the visualization of it is not alinged to the ground truth range, other wise its visualization is almost a single color. The implicit learned confidence also filters out the noise of monocular depth, especially in Figure 7.
@@
We provide additional visualizations of the intermediate results from global fusion in Figures 13, 14, 15, 16, 17, 18, 19, and 20. These visualizations illustrate the varying registration of monocular depth across individual pixels, particularly across different objects. Given that the monocular depth obtained from DepthAnything is scale ambiguous and does not represent absolute depth before registration, we do not align it with the ground truth range in visualization; otherwise, it would appear almost uniformly as a single color. The implicitly learned confidence also effectively filters out noise in the monocular depth as demonstrated in Figure 7.
Appendix F Future Work Discussion
@@
We present failure cases in Figures 21 and 22. In the first failure case, our method is confused by the glass door and glass window, where both the transparent surfaces and the behind scene are significant. Unlike simple transparent objects (e.g., a glass bottle), transparent scenes raise a new challenge for robotics, as they need to perceive both the transparent surface and the scene behind it. Failure to do so may cause robots to get stuck, for instance, when trying to reach an apple behind a glass window. If the robot perceives only the glass window, it will miss the apple entirely, while perceiving only the apple means the glass acts as an unrecognized and insurmountable barrier. Therefore, a novel representation for depth estimation is necessary to allow for multiple depths at a single pixel.
@@
Figure 13: The visualization of intermediate results. $itr$: the current iteration. $cz,cr,cq$: context used in GRU. $scale$: scale $0\sim 2$ represents resolution from high to low.
Figure 14: The visualization of intermediate results. $itr$: the current iteration. $cz,cr,cq$: context used in GRU. $scale$: scale $0\sim 2$ represents resolution from high to low.
Figure 15: The visualization of intermediate results. $itr$: the current iteration. $cz,cr,cq$: context used in GRU. $scale$: scale $0\sim 2$ represents resolution from high to low.
Figure 16: The visualization of intermediate results. $itr$: the current iteration. $cz,cr,cq$: context used in GRU. $scale$: scale $0\sim 2$ represents resolution from high to low.
Figure 17: The visualization of intermediate results. $itr$: the current iteration. $cz,cr,cq$: context used in GRU. $scale$: scale $0\sim 2$ represents resolution from high to low.
Figure 18: The visualization of intermediate results. $itr$: the current iteration. $cz,cr,cq$: context used in GRU. $scale$: scale $0\sim 2$ represents resolution from high to low.
Figure 19: The visualization of intermediate results. $itr$: the current iteration. $cz,cr,cq$: context used in GRU. $scale$: scale $0\sim 2$ represents resolution from high to low.
Figure 20: The visualization of intermediate results. $itr$: the current iteration. $cz,cr,cq$: context used in GRU. $scale$: scale $0\sim 2$ represents resolution from high to low.

### v2

@@ 4.4 Ablation Study and Analysis
The Effectiveness of Each Module. As shown in Table 5, the baseline model performs better without context features, indicating that monocular priors are susceptible to domain bias when data is limited. By incorporating less-biased monocular priors from a pre-trained large monocular network in the monocular encoder (ME), generalization performance is significantly improved, highlighting the importance of robust monocular priors in the wild world. Comparing the Baseline + ME with Baseline in Table 1, its performance becomes worse than Raft-Stereo, showing that it is easy to suffer from over-confidence when simply fusing monocular features with disparities during iterative disparity update. The iterative direct fusion method (IDF) fuses monocular depth and binocular disparity through direct concatenation and convolution at each iteration. Compared to this approach, our iterative local fusion (ILF) is more robust to noise in binocular disparity, resulting in superior performance. The post-fusion method (PF) fuses monocular depth with the optimized binocular disparity from the previous iteration without registration. Compared to this approach, our global fusion (GF) achieves better compatibility between monocular depth and binocular disparity, mitigating the noise caused by scale ambiguity during fusion. Our iterative local fusion and global fusion modules further enhance performance and improve model robustness with the monocular encoder. It is also noted that the Baseline time cost is 0.32s, while our model’s is 0.4s. Though it involves a VTF model, thanks to the elegant and controllable design, our model barely raises the time cost.
The Analysis of Iterative Local Fusion. We also analyze the specific configurations of iterative local fusion. As shown in Table 6, the fixed weights in the LBP-like operation have a slight impact on performance, with a kernel size of 3,5{3,5} providing optimal results. We also try to use convolutions with learnable weights to replace LBP-like convolutions. Comparing L(4), L(9-10) with L(6-7), we find that fixed-weight convolutions are more robust than learnable convolutions, and deeper learnable convolutions produce worse results. This is because limited data makes monocular-related learning unreliable for generalization, whereas manually designed convolutions incorporate prior knowledge and are less affected by data bias. Using a sigmoid function after LBP-like convolutions further improves overall performance. The amplitude parameter does not show a significant influence. We also visualize the local ordering map in Figure 3. The local ordering maps of predicted disparity gradually become similar to the result of monocular depth as the iteration increases. For more visualizations, please refer to our supplemental materials.
@@ 5 Conclusion
Acknowledgement This work was supported by the Shenzhen Science and Technology Program under Grant No. JCYJ20241202130548062, the Natural Science Foundation of Shenzhen under Grant No. JCYJ20230807142703006, the Natural Science Foundation of China (NSFC) under Grants No. 62176021 and No. 6217204, and the Key Research Platforms and Projects of the Guangdong Provincial Department of Education under Grant No.2023ZDZX1034.
Supplementary Material
[…]
[Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image]
Figure 5: The visualization of results on the Flicker1024 dataset.
Figure 6: The visualization of results. We use two kinds of colormap to visualize the disparity map.
The training and testing codes for all experiments, including the ablation study, are available in our project. For reproducibility, we strongly recommend referring to our project.
6 Visualization On Flicker1024
We present visualization results demonstrating the generalization capability of our model from the synthetic SceneFlow dataset to the real-world Flickr1024 dataset [47]. As shown in Figure 5, our model performs robustly across diverse scenarios, including large outdoor and indoor scenes, thin and small objects, strong lighting interference and low-light conditions, as well as challenging materials such as glass windows, walls, and bottles.
7 Intuition behind Monocular Depth Model
@@
Method
Additional
Data/Aug
Booster
@@
Method
Additional
Data/Aug
Booster
@@
8 More Results on Booster
We provide additional results on the Booster dataset across various material types. From class 0 to $3$, the materials become increasingly transparent and/or specular. As shown in Tables 8 and 9, our method outperforms state-of-the-art approaches on transparent and/or specular objects (classes $1$ to $3$), while achieving comparable results in normal regions (class 0). The normal regions of the Booster dataset mainly consist of regular objects, flat surfaces, or highly textured areas. Consequently, NerfStereo, which incorporates additional stereo data, performs particularly well in these regions. This indicates that stereo matching effectively captures fine-grained details, whereas monocular depth estimation excels in perceiving coarse shapes. As illustrated in Figures 7 and 8, binocular disparity provides greater detail compared to monocular depth. Our method disentangles monocular depth and binocular disparity, allowing the model to leverage both monocular and stereo data, and explore the fusion of monocular priors effectively.
9 Additional Training Data
We evaluate the scalability of our model by incorporating additional training data from the TranScene dataset [24], a synthetic dataset specifically designed for multi-label transparent scenes. In our experiments, we use labels with the largest disparity in transparent regions. It should be noted that, this time, our model is trained end-to-end using weights pretrained on the SceneFlow dataset, without using any multi-stage training strategy. As shown in Tables 10 and 11, incorporating the additional data leads to consistent performance improvements across all evaluated metrics, with particularly notable gains in transparent regions. Furthermore, our model’s performance on common scenes (e.g., non-transparent regions) not only remains stable but also shows slight improvement. These results highlight the scalability potential of our model when augmented with additional large data.
Metric ALL Trans NoTrans
Ours Ours+TranScene Ours Ours+TranScene Ours Ours+TranScene
EPE 2.26 1.24 7.93 5.67 1.52 0.75
RMSE 5.60 4.19 11.03 8.42 3.93 3.07
2px 11.02 7.91 59.83 46.78 6.98 4.77
3px 8.59 5.97 50.36 38.55 4.97 3.23
5px 6.60 4.52 38.44 28.65 3.64 2.29
6px 6.00 4.08 33.87 25.41 3.27 2.01
8px 5.35 3.44 27.56 21.30 2.89 1.59
Table 10: Generalization from the SceneFlow dataset to the Booster dataset in quarter resolution and balanced set. ‘All’, ‘Trans’, and ‘NonTrans’ represent all regions, transparent regions, and nontransparent regions, respectively.
Metric Class 0 Class 1 Class 2 Class 3
Ours Ours+TranScene Ours Ours+TranScene Ours Ours+TranScene Ours Ours+TranScene
EPE 0.79 0.75 1.53 1.40 5.32 1.62 7.93 5.67
RMSE 3.02 2.99 4.70 4.74 6.39 2.26 11.03 8.42
2px 5.90 5.15 12.67 9.17 23.34 13.51 59.83 46.78
3px 4.57 4.08 7.80 5.63 17.62 10.23 50.36 38.55
5px 3.17 3.00 4.88 3.80 13.50 7.40 38.44 28.65
6px 2.58 2.59 3.96 3.37 12.80 6.50 33.87 25.41
8px 1.45 1.73 3.14 2.86 12.15 4.93 27.56 21.30
Table 11: Generalization from the SceneFlow dataset to the Booster dataset in various regions.
[…]
@@ 9 Additional Training Data
[…]
@@
[…]
Table 13: The effectiveness of each module. Baseline: RAFTStereo, ME: our monocular encoder, ILF: iterative local fusion, GF: our global fusion. FE-DepthAnything: replacing the original feature extractor with DepthANything v2. FE-MASt3R: replacing the original feature extractor with MASt3R.
10 More Analysis about Memory
@@
11 More Visualization
@@
We provide additional visualizations of generalized stereo matching in Figures 9, 10, 11, 12, and 13. The visualizations span a variety of environments, ranging from open outdoor scenes (e.g., driving scenarios), to semi-open outdoor scenes (e.g., playgrounds), and to enclosed indoor scenes (e.g., rooms, tables). The results demonstrate that our method generalizes effectively to the wild world, achieving strong performance even when trained only on a limited amount of synthetic stereo data.
12 Ablation Study
12.1 More Analysis of Backbone
@@
In addition to replacing the context network with the pre-trained DepthAnything v2 [53], we also experimented with replacing the feature extractor for cost volume construction using DepthAnything v2 [53] and MASt3R [18, 45]. As shown in Table 13, the results become worse after replacing the feature extractor for cost volume construction with DepthAnything v2 or MASt3R. Moreover, a bug with the A40 GPU causes memory issues when converting the alternate correlation function from dot product to Euclidean distance during training. Therefore, the model with MASt3R was trained using the original correlation function with dot product, where additional learnable convolution layers are further used after MASt3R for feature extraction.
12.2 More Analysis of Iterative Local Fusion
@@
We provide additional visualizations of the intermediate results from the iterative local fusion process in Figures 14, 15, 16, 17, 18, 19, 20, and 21. As the iterations progress, the ordering maps generated from binocular disparity gradually become smoother. The convolution layers learn the differences between ordering maps generated from binocular disparity and monocular depth, allowing the guidance to focus more effectively on non-smooth regions, thereby significantly affecting disparity update.
12.3 More Analysis of Components in Global Fusion
@@
We present more visualization for the intermediate results of global fusion in Figure 14, 15, 16, 17, 18, 19, 20, and 21. The visualization shows that the registration of monocular depth is different for each pixel, particularly on different objects. Since the monocular depth from DepthAnything is scale ambiguity but not absolute depth before registration, the visualization of it is not alinged to the ground truth range, other wise its visualization is almost a single color. The implicit learned confidence also filters out the noise of monocular depth, especially in Figure 8.
@@
We provide additional visualizations of the intermediate results from global fusion in Figures 14, 15, 16, 17, 18, 19, 20, and 21. These visualizations illustrate the varying registration of monocular depth across individual pixels, particularly across different objects. Given that the monocular depth obtained from DepthAnything is scale ambiguous and does not represent absolute depth before registration, we do not align it with the ground truth range in visualization; otherwise, it would appear almost uniformly as a single color. The implicitly learned confidence also effectively filters out noise in the monocular depth as demonstrated in Figure 8.
13 Future Work Discussion
@@
We present failure cases in Figures 22 and 23. In the first failure case, our method is confused by the glass door and glass window, where both the transparent surfaces and the behind scene are significant. Unlike simple transparent objects (e.g., a glass bottle), transparent scenes raise a new challenge for robotics, as they need to perceive both the transparent surface and the scene behind it. Failure to do so may cause robots to get stuck, for instance, when trying to reach an apple behind a glass window. If the robot perceives only the glass window, it will miss the apple entirely, while perceiving only the apple means the glass acts as an unrecognized and insurmountable barrier. Therefore, a novel representation for depth estimation is necessary to allow for multiple depths at a single pixel.
@@
Figure 14: The visualization of intermediate results. itritr: the current iteration. $cz,cr,cq$: context used in GRU. scalescale: scale $0\sim 2$ represents resolution from high to low.
Figure 15: The visualization of intermediate results. itritr: the current iteration. $cz,cr,cq$: context used in GRU. scalescale: scale $0\sim 2$ represents resolution from high to low.
Figure 16: The visualization of intermediate results. itritr: the current iteration. $cz,cr,cq$: context used in GRU. scalescale: scale $0\sim 2$ represents resolution from high to low.
Figure 17: The visualization of intermediate results. itritr: the current iteration. $cz,cr,cq$: context used in GRU. scalescale: scale $0\sim 2$ represents resolution from high to low.
Figure 18: The visualization of intermediate results. itritr: the current iteration. $cz,cr,cq$: context used in GRU. scalescale: scale $0\sim 2$ represents resolution from high to low.
Figure 19: The visualization of intermediate results. itritr: the current iteration. $cz,cr,cq$: context used in GRU. scalescale: scale $0\sim 2$ represents resolution from high to low.
Figure 20: The visualization of intermediate results. itritr: the current iteration. $cz,cr,cq$: context used in GRU. scalescale: scale $0\sim 2$ represents resolution from high to low.
Figure 21: The visualization of intermediate results. itritr: the current iteration. $cz,cr,cq$: context used in GRU. scalescale: scale $0\sim 2$ represents resolution from high to low.